OpenStudy (anonymous):

solve for all values of x for which the expression is defined, and check: 1/x-1=x-1/4 express your answers in ascending order: x=_ or x =__ please help!

OpenStudy (anonymous):

\[\frac{1}{x-1}=\frac{x-1}{4}\]?

OpenStudy (anonymous):

that is my guess anyway if so then cross multiply and get \[(x-1)^2=4\] so \(x-1=2\) or \(x-1=-2\)